OpenStudy (anonymous):

(x-4)^2=(*9x*)^2 stars=radical sign

OpenStudy (anonymous):

\[(\sqrt{a} )^2=a\]

OpenStudy (anonymous):

but both sides got the square

OpenStudy (anonymous):

(x-4)^2 foil it

OpenStudy (anonymous):

are u sure? what about the other side

OpenStudy (anonymous):

are side would be just 9x

OpenStudy (anonymous):

ok so the original equation was (x-4)=9x but the 9x was in a radical sign does the foil thing apply?